3x2 − 4y2 + 6x + 24y − 135 = 0
1
Expert's answer
2020-06-08T17:27:40-0400

"3x^2 \u2212 4y^2 + 6x + 24y \u2212 135 = 0"

"3(x^2 +2x+1-1)\u22124(y^2 - 6y+9-9) \u2212 135 = 0"

"3(x-1)^2\u22124(y - 3)^2 \u2212 168 = 0"

"(x-1)^2\/56-(y-3)^2\/42=1"

Thus, the conic represented by the given equation is a hyperbola whose focus is at (1,3).